The antifungal drug anidulafungin is administered by IV infusion. The maximum infusion rate is achieved when 260 ml of anidulafungin solution, which has a concentration of 0.77 mg/ml, are delivered in 3 hours. What is the maximum infusion rate in milligrams per minute?

A loading dose of voriconazole consists of 5 mg for every kilogram of body mass. This dose is administered as two IV infusions separated by 12 hours. The concentration of the solution is 10 mg/ml. If a patient weighs 132 pounds, how many milliliters of voriconazole are administered in a day?

Valacyclovir is an antiviral drug that interferes with the replication process of a virus. For the treatment of Herpes zoster, the typical dose consists of 1000 mg administered three times a day over a period of 7 days. How many 1000 mg tablets of valacyclovir are needed?

An IV infusion of ganciclovir is administered to a patient whose weight is 187 pounds. A dose of ganciclovir is administered at a rate of 5 mg/hr for every kilogram of body mass. Two 1-hour infusions are given each day for a minimum of 14 days. If the concentration of ganciclovir is 50 mg/ml, how many milliliters are required for a two-week regimen of ganciclovir?
Answer #1

1. The maximum infusion rate is achieved when 260 ml of anidulafungin solution, which has a concentration of 0.77 mg/ml, are delivered in 3 hours.

Means 0.77mg in 1 ml

Or 1 ml contains 0.77 mg

Then 260 ml will contain=0.77X260=200.2 mg

Then infusion rate in milligrams per minute= 200/ (3X60) = 1.11mg/min

2.Loading dose = 5 mg/kg

Patient weighs 132 pounds=60 kg.

Total dose as per weight will be required=60X 5=300mg

Available concent. = 10 mg/ml.

Required dose in ml=?

10mg in 1 ml

1mg in=1/10 ml

300mg will be available in/ supplied by the=(1/10)X300=30 ml

3.For the treatment of Herpes zoster, the typical dose consists of 1000 mg administered three times a day over a period of 7 days.

Per day amount =3000mg

For 7 days=7x3000=21000mg

How many 1000 mg tablets of valacyclovir are needed?
1000mg by 1 tablet then 21000mg from 21 tablets will be required

4. A dose of ganciclovir is administered at a rate of 5 mg/hr for every kilogram of body mass. Patient weight= 187 pounds= 85kg

Total required dose as the weight is concerened=85X5=425mg/hr

Available concentration of ganciclovir is= 50 mg/ml

425mg=425/50ml=8.5 ml

Required infusion in ml/hr=8.5ml/hr

Two 1-hour infusions are given each day for a minimum of 14 days.

Then per day required ml=8.5X 2=17ml/day

How many milliliters are required for a two-week regimen of ganciclovir?

17ml/day

Then for two week regimen requiredml=17X14=238ml